ganeshie8 (ganeshie8):

quick question : if we are given f(f(x)) = f(x), f:X->X is it okay to conclude f(x) = x ?

OpenStudy (kainui):

f(x)=C or f(x)=0 seem to be possibilities as well.

OpenStudy (kainui):

Or actually the math floor, math ceiling, and absolute value of x functions also have this property.

OpenStudy (kainui):

http://en.wikipedia.org/wiki/Idempotence

OpenStudy (vishweshshrimali5):

Yes you can

OpenStudy (vishweshshrimali5):

If I substitute f(x) = y then : \[\large{f(y) = y}\] Which is similar to \[\large{f(x) = x}\]

OpenStudy (vishweshshrimali5):

Provided that domain and range are same ;) :P

ganeshie8 (ganeshie8):

f(x) = c f(f(x)) = f(c) = c oh my bad yes :o whats wrong with below argumennt : \[f(f(x)) = f(x)\] taking inverse both sides gives \[f(x) = x\] right ?

ganeshie8 (ganeshie8):

yeah

OpenStudy (vishweshshrimali5):

Well both are correct but ensure that domain and range satisfy the conditions

OpenStudy (kainui):

That's a cool way of doing it, I was wondering how you came to that answer. Nice. =)

OpenStudy (vishweshshrimali5):

\(\large{fof(x)}\) and other composite functions are very stubborn while dealt with domain and range

OpenStudy (vishweshshrimali5):

@ganeshie8 Solving the combinatorics question ? ;)

ganeshie8 (ganeshie8):

yes trying lol so f(x)=x doesn't cover all the functions for which f(f(x)) = x is it ? cuz f(x) =c is certainly an outlier

OpenStudy (vishweshshrimali5):

Yes inverse of function will exist only if certain domain and range conditions are fulfilled. It would be best if you use your question along with the restrictions assumed to be true

OpenStudy (vishweshshrimali5):

Finally you can find out the total number of possibilities if you need :)

ganeshie8 (ganeshie8):

Oh yes inverse need not exist for f, thats another part ive missed this info is given in the question : `f : X->X, X has "n" elements`

ganeshie8 (ganeshie8):

that wiki link has the formula for total # of functions possible : \[\large \sum \limits_{k=0}^n\binom{n}{k}k^{n-k}\]

OpenStudy (vishweshshrimali5):

Yup correct formula :) But would it help ?

OpenStudy (vishweshshrimali5):

If that info is given then inverse will exist

ganeshie8 (ganeshie8):

\[\large \sum \limits_{k=0}^n\binom{n}{k} = 2^n\] refers to powerset, and i think it represents the total number of functions for which f(x) = x

ganeshie8 (ganeshie8):

* refers to powerset, and i think it represents the total number of functions of form : f(x) = x

ganeshie8 (ganeshie8):

\(\large k^{n-k}\) is somehow accounting for other functions :o

OpenStudy (vishweshshrimali5):

One more thing: for inverse to exist the function must be one-one and onto. Ensure that this condition is also fulfilled

OpenStudy (vishweshshrimali5):

Or you can just use a formula to find out the number of such functions :)

OpenStudy (vishweshshrimali5):

Then state that out of these possible functions one is f(x) Not sure if it is going to help... I am not that good at combinatorics :D

ganeshie8 (ganeshie8):

so if i understand it correctly : f(x) = x essentially represents all the invertible funcitons for which f(f(x)) = f(x) is true and there could be other functions without inverses of form f(f(x)) = f(x) ?

OpenStudy (vishweshshrimali5):

I am not sure... the problem is if I think that fof(x) = f(x) is true then it appears to me that even stating that f(x) belongs to the domain of x, then it should be invertible because in any other case if would not be true *shrug*

OpenStudy (vishweshshrimali5):

Because the equation probably automatically ensures that the function is invertible. If we think along this path then it should imply that function is both one one and onto. While this can provide some more results regarding domain and range of the function, it probably would not be able to solve the question

ganeshie8 (ganeshie8):

I see... in that case we can safely take inverse both sides and conclude f(x)=x right ?

OpenStudy (vishweshshrimali5):

Yeah I think so

ganeshie8 (ganeshie8):

then that formula in wiki must be representing all the funcitons of form : f(x) = x

OpenStudy (vishweshshrimali5):

Probably

ganeshie8 (ganeshie8):

Oh wait, what about the outlier : f(x) = c f(x) = x doesn't capture this function

OpenStudy (vishweshshrimali5):

This is the main problem that I too am facing

ganeshie8 (ganeshie8):

then good :) we're one same page xD

OpenStudy (vishweshshrimali5):

:D

OpenStudy (vishweshshrimali5):

I was on the same page from yesterday evening IST

OpenStudy (vishweshshrimali5):

:P

ganeshie8 (ganeshie8):

haha same here lol that wiki link looks useful, i am going thru idempotence stuff

OpenStudy (vishweshshrimali5):

Okay :) Hope that you find something helpful from the link